OpenStudy (anonymous):

The Cartesian coordinates of a point in the xy-plane are (x, y) = (-3.06, -2.24) m. Find the polar coordinates of this point. r = θ =

OpenStudy (badhi):

its obvious that \(\theta\) should be between 180 and 270 degrees since both x and y coordinates are negative

OpenStudy (anonymous):

BAdhi is right. \[\theta \] should be in between 180 and 270 degrees. r is \[\sqrt{?} \](xx+yy) . and then . \[\theta \] =\[\tan^{-1} \] y/x . :)

OpenStudy (anonymous):

lol sorry im new here so dont know how to write those equations properly. r is the square root of (x^2 + y^2) and theeta is tan inverse (y/x)

OpenStudy (irishboy123):

if in doubt, draw :-)